Divide $80 among three people so that the second will have twice as much as the first, and the third will have $5 less than the second.(x=17)
Answer by mananth(16949)

let first have x
second has 2x
third has (2x-5)

x+2x+2x-5=80
5x-5=80
5x=85
/5
x=17
17, 34,29